For steady level cruise the engine thrust must equal the drag, which is given by the aircraft weight divided by the lift drag ratio of the aircraft. There must be additional engine thrust if the aircraft is to climb. If at 35,000 ft the flight Mach number is 0.78 and the rate of climb is 500 ft/min, find the angle of climb. If L/D = 21.6, find the proportional increment in net thrust needed to achieve this climb rate.

(ans: 23.7%)
Answer #1

Thrust= W / (L/D)

w=L/ cos\theta

climb angle sin(theta)=T/W-1/(L/D)